Events $E$ and $F$ are such that $P(\text{not } E \text{ and not } F) = 0.25$. State whether $E$ and $F$ are mutually exclusive.

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(N/A) It is given that $P(\text{not } E \text{ or not } F) = 0.25$.
By De Morgan's Law,this is equivalent to $P((E \cap F)') = 0.25$.
We know that $P(E \cap F) = 1 - P((E \cap F)')$.
Substituting the value,we get $P(E \cap F) = 1 - 0.25 = 0.75$.
Since $P(E \cap F) = 0.75 \neq 0$,the events $E$ and $F$ are not mutually exclusive.

Explore More

Similar Questions

$A$ bag contains $6$ white,$5$ black,and $4$ red balls. Find the probability of getting a white or a black ball when one ball is drawn from the bag.

When two dice are rolled,what is the probability that the sum of the numbers on the faces is a multiple of $4$?

$A$ die is formed so that the probability of getting a number $i$ when it is rolled is proportional to $i$ $(i=1, 2, 3, 4, 5, 6)$. The probability of getting an odd number on the die when it is rolled is

$P, Q$ and $R$ try to hit the same target one after the other. If their probabilities of hitting the target are $\frac{2}{3}, \frac{3}{5}, \frac{5}{7}$ respectively,then the probability that the target is hit by $P$ or $Q$ but not by $R$ is

If two fair dice are rolled,then the probability that the sum of the numbers on the upper faces is at least $9$,is

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo